OpenStudy (anonymous):

Which of the following is not an element of the following set: A: {positive integers divisible by 3} 3 21 14 30

OpenStudy (anonymous):

3 / 3 = 1 21 / 3 = 7 14 / 3 = 4.666666666 30 / 3 = 10 u decide :)